### v1

@@ 5.2 Limitations across Proposed APIs
In addition to gaps in the MPI standard that increase the difficulty of supporting GPU triggering, there are also recurring limitations across the various APIs that have been proposed. We highlight these gaps below to encourage efforts to address them consistently in new proposals:
  • Clear concurrency model—If the host execution context and accelerator context both initiate operations, the ordering and concurrency is ill-defined in many APIS; any (left to the vagueness of MPI), or are silent. MPICH stream triggering is the only one that provides a clear, specific concurrent model for both host- and accelerator-initiated communication.
  • Matching—There is no explicit support for matching of send and receive buffers so that this occurs only once, and receive queue matching overheads are eliminated.
  • GPU completion support with partitioning—If something is initiated on a GPU, there is no way to know if something is completed on the send side. This makes it difficult or impossible to write programs that do not need to synchronizae with the CPU to interleave communication and computation.
  • Collective communication support—This is broadly absent and needed to support real applications that avoid unnecessary synchronization with the CPU. In particular, no current stream triggering approach supports full stream-triggered collective communication.
  • GPU/NIC Progress—Lack of this feature is a gap for any performant implementation. Out of the eight APIs considered only two (partitioned communication augmented with Pbuf_prepare and HPE one-sided communication) support GPU/NIC progress.
@@ 6 Conclusion
This is the first comprehensive survey of all eight proposals known to the authors that address stream/graph- and kernel-triggered communication abstractions for MPI, whose principal purpose is to improve the performance of communication when GPU kernels create or consume data for transfer through MPI operations. As stated at the outset, GPU-enhanced architectures are now dominant in HPC systems, but message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches that lower such costs.
Here, we compared and contrasted stream/graph- and kernel-triggered MPI communication abstractions. Various proposals with distinct potential APIs for stream- and/or kernel-triggering have been proffered as designs and/or prototypes, which span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, and explicit MPI stream/queue objects. Notably, some of these proposals either strengthen or weaken the semantics of MPI operations. We enumerated eight different proposals and provided a taxonomy of these proposals and their operations against twelve categories of features and semantics.
We described the design space in which these abstractions reside, their implicit or explicit use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discussed their potential for added performance, how usable these abstractions are, and where functional and/or semantic gaps exist. Designs breaking backward compatibility with MPI were also duly noted (and cautioned). We also presented code examples of how some of these operate on the ubiquitous ping-pong use case (some utilize CUDA and others ROCM APIs).
The next steps to advance performance-portable programming with accelerator-based MPI+X programming should start with robust, inclusive conversations and design bake-offs among the various proposers, implementers, and users to identify a coherent path forward within the MPI community, inclusive of application developers. One or more proposed standard APIs should be competed and down-selected for MPI-5 (or an earlier MPI-4.x release). Significantly, the existing proposals/prototypes cited in this paper may not include all hardware vendors or systems currently available and certainly there may be other solutions forthcoming or that are not yet public (e.g., for FPGA-based accelerators or other GPU accelerator products). Yet, fitting additional and/or forthcoming proposals, prototypes, and designs into our taxonomy will provide a coherent basis to compare and contrast these with the eight presented here.
@@ 6.0.1 Acknowledgements
This work was performed with partial support from the National Science Foundation under Grants Nos.  OAC-2103510, CCF-2405142 and CCF-2412182, the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966, and Tennessee Technological University.
Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, or the U.S. Department of Energy’s National Nuclear Security Administration.

### v3

@@ 5.2 Limitations across Proposed APIs
In addition to gaps in the MPI standard that increase the difficulty of supporting GPU triggering, there are also recurring limitations across the various APIs that have been proposed. We highlight these gaps below to encourage efforts to address them consistently in new and revised proposals:
  • Clear concurrency model—If the host execution context and accelerator context both initiate or start operations, the ordering and concurrency is ill-defined in many APIs, are left to the vagueness of MPI, or are silent. MPICH stream triggering is the only one that provides a clear, specific concurrency model for both host- and accelerator-initiated communication.
  • Matching—There is no explicit support for matching of send/receive buffers so that this occurs only once, and receive matching overheads are eliminated.
  • GPU completion support with partitioning—If something is started on a GPU, there is no way to know if something is completed on the send side. This makes it difficult or impossible to write programs without need to synchronize with the CPU to interleave communication and computation.
  • Collective communication support—This is broadly absent yet needed to support real applications that avoid unnecessary synchronization with the CPU. In particular, no current MPI stream triggering approach supports stream triggering of the full range of MPI collective communication operations.
  • GPU/NIC Progress—Lack of this feature is a gap for any performant implementation. Out of the nine APIs considered only three (partitioned communication augmented with Pbuf_prepare, HPE one-sided communication, and Intel GPU-initiated) support GPU/NIC progress.
@@ 6 Conclusion
Message-passing communication involving GPUs with MPI has proven to be both complex and expensive, motivating new approaches. This is the first comprehensive survey of all nine proposals known to the authors that address stream/
graph-, kernel-triggered, and GPU-initiated communication abstractions for
MPI, whose principal purpose is to improve the performance of communication when GPU kernels create or consume data for transfer through MPI operations.
Here, we compared and contrasted stream/graph- and kernel-triggered MPI communication abstractions. Various proposals with distinct potential APIs for stream- and/or kernel-triggering have been proffered as designs and/or prototypes, which span various GPU architectures and approaches, including MPI-4 partitioned point-to-point communication, stream communicators, explicit MPI stream/queue objects, and notified-RMA. Notably, some of these proposals either strengthen or weaken the semantics of MPI operations. We enumerated nine different proposals and provided a taxonomy of these proposals and their operations against twelve categories of features and semantics.
We described the design space in which these abstractions reside, their use of stream and other non-MPI abstractions, their relationship to partitioned and persistent operations, and discussed their potential for added performance, how usable these are, and where functional and/or semantic gaps exist. Designs breaking backward compatibility with MPI were also duly noted (and cautioned). We also presented code examples of how some of these operate on the ubiquitous ping-pong use case (variously utilizing CUDA, ROCM, and Intel APIs).
The next steps to advance performance-portable programming with accelerator-based MPI+X programming should start with robust, inclusive conversations and design bake-offs among the various proposers, implementers, and users to identify a coherent path forward within the MPI community, inclusive of application developers. One or more proposed standard APIs would be competed then down-selected for MPI-5. Existing proposals/prototypes cited in this paper may not include all hardware vendors or systems currently available and certainly there may be other solutions forthcoming or that are not yet public (e.g., for FPGAs or other GPU accelerator products). Yet, fitting additional and/or forthcoming proposals, prototypes, and designs into our taxonomy will provide a coherent basis to compare and contrast these with the nine presented here.
@@ 6.0.1 Acknowledgements
The authors wish to acknowledge Dr. Daniel J. Holmes of Intel Corporation for providing information, examples, and references on the Intel GPU triggering API for this paper. The authors also wish to acknowledge Dr. Hui Zhou of Argonne National Laboratories for providing additional information and clarifications on the MPICH stream triggering API.
This work was performed with partial support from the National Science Foundation under Grants Nos.  OAC-2103510, CCF-2405142 and CCF-2412182, the U.S. Department of Energy’s National Nuclear Security Administration (NNSA) under the Predictive Science Academic Alliance Program (PSAAP-III), Award DE-NA0003966, and Tennessee Technological University. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation, or the U.S. Department of Energy’s National Nuclear Security Administration.